OpenStudy (anonymous):

A rectangle has an area of 21.42 cm2. When both the length and width of the rectangle are increased by 1.50 cm, the area of the rectangle becomes 38.22 cm2. Calculate the length of the longer of the two sides of the initial rectangle.

OpenStudy (amistre64):

would you agree that length times width is area?

OpenStudy (amistre64):

l w = k (l+n) (w+n) = c let, say l=k/w and sub in (k/w+n) (w+n) = c multiplying by w we get (k+wn) (w+n) = wc which is just a quadratic in w to determine

OpenStudy (amistre64):

kw + kn + nw^2 + n^2 w = wc n w^2 + (n^2 + k - c) w + kn = 0 for the given values of n,k,c from the problem

OpenStudy (amistre64):

the 2 values you get for w, will actually define the 2 required sides, pick the larger for the solution

OpenStudy (anonymous):

what does k stand for?

OpenStudy (amistre64):

the area of a rectangle, before adding something to the sides ... lw = k is simpler to work with instead of writing in all the numbers to get in the way

OpenStudy (anonymous):

im still kinda confused, do you think you could explain it some more?

OpenStudy (anonymous):

is c the area after adding? what is n?

OpenStudy (amistre64):

c is the area after adding, n is what they added yes

OpenStudy (amistre64):

\[(1.5) w^2 + ((1.5)^2 + (21.42) - (38.22)) w + (21.42)(1.5) = 0\] still messy, might have been better to get teh quadratic formula first then inserted values :)

OpenStudy (amistre64):

\[\underbrace{n}_a~ w^2 +\underbrace{ (n^2 + k - c)}_b~ w + \underbrace{kn}_c = 0\] \[w=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] still a mess but at least we dont have to fight the numbers till the end

OpenStudy (amistre64):

6.3 * 3.4 = 21.42 and (6.3+1.5) * (3.4+1.5) = 38.22 as required

OpenStudy (anonymous):

thank you!

OpenStudy (amistre64):

who ever assigned this to you .. you need to slap upside the head :) but at least you tried to participate and i hope my method proved insightful at least.